### v1

@@ Abstract
This paper presents JavisGPT, the first unified multimodal large language model (MLLM) for Joint Audio-Video (JAV) comprehension and generation. JavisGPT adopts a concise encoder–LLM–decoder architecture, featuring a SyncFusion module for spatio-temporal audio-video fusion and synchrony-aware learnable queries to bridge a pretrained JAV-DiT generator. This design enables temporally coherent video-audio understanding and generation from multimodal instructions. We design an effective three-stage training pipeline consisting of multimodal pretraining, audio-video fine-tuning, and large-scale instruction-tuning, to progressively build multimodal comprehension and generation from existing vision-language models. To support this, we further construct JavisInst-Omni, a high-quality instruction dataset with over 200K GPT-4o-curated audio-video-text dialogues that span diverse and multi-level comprehension and generation scenarios. Extensive experiments on JAV comprehension and generation benchmarks show that JavisGPT outperforms existing MLLMs, particularly in complex and temporally synchronized settings.
@@ 1 Introduction
Joint understanding and generation of multiple modalities has recently sparked increasing interest in the community [tang2023codi, wu2024towards, fei2024vitron, tang2024codi2, xu2025qwen25omni, microsoft2025phi4minitechnicalreportcompact], owing to its broader applicability compared to unimodal systems. However, current efforts predominantly focus on image-text frameworks [zhou2024transfusion, xie2024show, team2024chameleon, pan2025metaquery, chen2025januspro], or treat video and audio understanding and generation as separate modalities [zhan2024anygpt, wu2024next, lu2024unifiedio2]. In contrast, truly unified modeling of synchronized sounding videos—which requires precise temporal alignment between audio and visual streams for both comprehension and generation—remains significantly underexplored. In this paper, we aim to bridge this gap by designing an MLLM capable of jointly understanding and generating videos with synchronized audio (i.e., sounding video). This is a critical yet practical problem with broad real-world applications [xu2025qwen25omni, polyak2024movie, kong2024hunyuanvideo], such as dynamic avatar chatbots, sounding movie curation, and music-driven video analysis.
@@
We introduce JavisGPT, a bespoke MLLM for high-quality, unified comprehension and generation in joint audio-video (JAV) tasks. As its core, JavisGPT is a holistic system that accepts as input separate audio and videos, synchronized sounding videos, and user text prompts. After comprehending and reasoning over these input signals, it produces either textual responses and/or synchronized sounding videos. Our architecture supports a wide range of JAV scenarios, including question-answering in comprehending sounding audios, generation of sounding audios following user instructions, upon in-context reasoning, and interleaved text-audio-video generation in multi-turn user dialogues (ref. Fig.˜1). As a comprehensive and state-of-the-art (SOTA) system, we highlight the design of JavisGPT below (with detailed discussions with related works provided in Sec.˜2).
Model architecture. JavisGPT adopts an encoder-LLM-decoder architecture (ref. Fig.˜2), with Qwen2.5 [yang2024qwen25] as the LLM backbone. The visual encoder is inherited from Qwen2.5-VL [bai2025qwen25vl], and the audio encoder is based on BEATs [chen2022beats]. Audio and video features, along with user prompts and learnable JavisQuery tokens, are passed to the LLM. To enable fine-grained spatiotemporal alignment, we propose a dedicated SyncFusion module that fuses audio and video representations into synchronized SyncAV tokens for unified comprehension. At the output stage, the LLM generates textual responses along with JavisCond tokens, which encode contextual semantics and serve as conditioning inputs for a pretrained JAV-DiT generator [liu2025javisdit]. We choose JavisDiT for its generation quality and flexibility, and incorporate hierarchical JavisQueries to provide spatiotemporal priors, further enhancing synchronization in audio-video generation.
@@
We compare JavisGPT with existing JAV-like LLMs on various JAV benchmarks, where JavisGPT yields SOTA performance across all scenarios. Moreover, for audio-video joint generation, JavisGPT also significantly outperforms existing methods, producing high-quality and diverse sounding videos. We reiterate our main points below:
@@ 3.1 Synchrony-aware Audio-Video Comprehension
Synchronized audio-video fusion. Unlike previous MLLMs that straightforwardly concatenate video and audio features together [cheng2024videollama2, sun2024videosalmonn] or simply interleave audio-video frames [tang2024videosalmonn2, xu2025qwen25omni], JavisGPT introduces a SyncFusion module $\psi^{av}$ to explicitly model the spatio-temporal synchrony in sounding videos. As illustrated in Fig.˜3, we first temporally align the audio and video streams by evenly dividing the audio representations to match the number of video frames, transforming the audio features from $\mathbf{A}\in\mathbb{R}^{T_{a}\times M\times C}$ to $\mathbf{A}^{\prime}\in\mathbb{R}^{T_{v}\times(r\times M)\times C}$, where $r=\lceil T_{a}/T_{v}\rceil$, with zero-paddings applied if needed. Then, a shared cross-attention module is employed to inject audio cues into the frame-wise video representations, followed by a two-layer MLP to generate SyncAV representations. The length of SyncAV representations matches that of the vanilla extracted video representations, ensuring compatibility with the MRoPE positional encoding used in the Qwen2.5-VL [bai2025qwen25vl]. Each resulting token $e_{i,j}^{t_{v}}\in\mathbb{R}^{C}$ encodes a localized audio-visual event occurring at the $i$-th row, $j$-th column visual patch of frame $t_{v}$. By integrating audio semantics into visual features, the SyncFusion module enhances representational expressiveness and achieves precise spatio-temporal synchronization—crucial for downstream tasks requiring fine-grained audio-video understanding.
@@ 3.2 Spatio-Temporally Synchronized Audio-Video Generation
Synchrony-enhanced audio-video generation. To further strengthen the spatio-temporal synchrony of generated sounding videos, we follow liu2025javisdit to additionally design $N^{\prime}$ learnable query tokens $Q^{s}$ to estimate spatio-temporal prior (ST-prior) embeddings $\hat{\mathbf{s}}$ for auxiliary conditions. As shown in Fig.˜4, the semantic condition $\hat{\mathbf{c}}$ a coarse description of the overall audio-visual event (e.g., ’a car starts its engine and leaves the screen’), while the ST-prior condition $\hat{\mathbf{s}}$ captures fine-grained synchrony. Specifically, the spatial prior governs where events occur and disappear (e.g., ’the car is in the top-left corner of screen’), whereas the temporal prior determines when they begin and end (e.g., ’sound starts at 2s, exits at 7s, fades at 9s’). To compute $\hat{\mathbf{s}}$, we employ a separate MLP projector $\phi_{\theta}^{s}$ to map the final-layer representations of the prior query tokens into the ST-prior condition space, aligning them with the ST-prior encoder used in JavisDiT [liu2025javisdit]. Finally, we combine the alignment loss between the predicted ST-prior embeddings $\hat{\mathbf{s}}$ and the ground-truth ST-priors $\mathbf{s}$ with the denoising objective, jointly optimizing both to strengthen the audio-video synchrony:
@@ 4 Training and Datasets
Table 1: Summary of the training pipeline. ’$M$-PT’ denotes multimodal pre-training, ’AV-FT’ refers to audio-video fine-tuning, and ’$M$-InstTune’ means multimodal instruction-tuning.
@@ 5 Experiments
This section presents the experimental results to evaluate the effectiveness of our proposed JavisGPT. For a fair comparison, we train JavisGPT following the three-stage pipeline described in Sec.˜4, with each stage running for one epoch. We bring out zero-shot evaluations on diverse downstream tasks covering audio-video comprehension and generation using their official evaluation protocols. Additional implementation and training details can be found in Appendix˜B.
@@ 5.1 Comparison with the State of the Art
Multimodal comprehension. As shown in Tab.˜2, our method achieves performance on par with SOTA models on video/audio understanding, demonstrating strong unimodal capabilities. It maintains visual comprehension strength of Qwen2.5-VL [bai2025qwen25vl] backbone (70.2 vs. 70.5 on Perception [patraucean2023perception]) and even slightly outperforms Qwen2.5-Omni [xu2025qwen25omni] on MVBench [li2024mvbench] (video) and TUT2017 [mesaros2017tut] (audio). For audio-video synchronized understanding, JavisGPT surpasses both Qwen2.5-Omni and VideoLLaMA2.1 [cheng2024videollama2] on AVQA [yang2022avqa] and MU-AVQA [liu2024musicavqa], achieving top results on both benchmarks. We attribute this to the spatio-temporal alignment capability of the proposed SyncFusion and our synchrony-aware training strategy. Notably, those results are obtained with fewer training samples (1.5M in total), highlighting its strong data efficiency.
@@
Interleaved comprehension and generation. As no existing benchmark directly measures joint understanding and generation of sounding videos, we fill this gap by carefully curating a new evaluation set comprising 100 multi-turn QA dialogues covering four types of interleaved understanding-generation tasks. We employ three unbiased volunteers to conduct human evaluation over five dimensions, covering both textual and audio-video outputs. Full setup and detailed analysis are provided in Sec.˜E.1 The average results are reported in Fig.˜6, where our method consistently outperforms the two competing methods with joint comprehension-generation capabilities by a large margin. Notably, we observe (1) NExT-GPT [wu2024next] frequently refuses to respond or generates noisy audio-video outputs, severely impacting instruction-following and generation quality; (2) UnifiedIO-2 [lu2024unifiedio2] is inferior to context reasoning and proactive thinking, failing to connect user history (e.g., preferences) with current instructions.
@@ 5.2 Further Analysis
Joint training improves both understanding and generation. Sec.˜5.2 compares the performance of comprehension and generation tasks under separate vs. joint training, with each score averaged over the corresponding test sets (e.g., ’V-Und’ is the average accuracy over ActivityNet [caba2015activitynet], Perception [patraucean2023perception], and MVBench [li2024mvbench]). Results show that joint training consistently outperforms separate training in both understanding and generation tasks, indicating a mutual enhancement effect. The improvement is more substantial on the generation side.
To better understand this effect, Sec.˜5.2 presents the evolution of generation metrics across three dimensions. As the number of generation training samples increases, the inclusion of comprehension data leads to steady gains in both AV-Quality and Text-Consistency. It suggests that semantic understanding of AV content enhances the quality of condition embeddings used for generation—a trend aligned with findings in MetaQueries [pan2025metaquery] for image tasks. However, AV synchrony shows only marginal improvement (+0.04), possibly indicating the need for a more unified mechanism bridging comprehension and generation. We further discuss this in Appendix˜A.
Case study: instruction-followed and synchrony-aware comprehension and generation. Fig.˜9 presents a qualitative comparison with UnifiedIO-2 [lu2024unifiedio2] and NExT-GPT [wu2024next] in a two-round dialogue setting. In the first round, the models are asked to identify the source of a gunfire sound. UnifiedIO-2 gives a general response (’a semi-automatic weapon’), while NExT-GPT hallucinates an explosion of a flower. In contrast, our model accurately grounds the sound in a video game scene, showing better spatio-temporal grounding and scene understanding. In the second round, the models are asked to generate a gunfire-related video. UnifiedIO-2 produces irrelevant key frames, NExT-GPT fails to generate, while our JavisGPT successfully synthesizes coherent and synchronized audio-visual content. More examples are presented in Appendix˜F.
@@ 6 Conclusion and Discussion
In this work, we propose JavisGPT, a unified MLLM for synchronized AV comprehension and generation. Built upon an encoder-LLM-decoder architecture, JavisGPT introduces a SyncFusion module for spatio-temporal alignment and integrates a pretrained JavisDiT generator via hierarchical condition embedding. Moreover, we establish a three-stage training pipeline tailored for synchronized generation and curate a large-scale JavisInst-Omni dataset for complex audio-video-text interactions. Extensive experiments demonstrate state-of-the-art performance on both unimodal and synchronized JAV tasks, with superior generation quality, consistency, and synchrony. We believe JavisGPT offers a strong foundation for future research in synchronized media generation.
Discussion 1: the extended capabilities for sounding video comprehension and generation. In the future, JavisGPT can be further extended to achieve higher fidelity in several directions: (1) Support for speech input and output [xu2025qwen25omni], which can be enabled by replacing the current BEATs [chen2022beats] tokenizer with more advanced alternatives such as Whisper [radford2023whisper] or WavTokenizer [ji2024wavtokenizer]; (2) Fine-grained controllable audio-video generation and editing, which requires the downstream JAV-DiT to be conditioned on additional control signals for precise generation [jiang2025vace]; (3) Complex instruction-based generation, which involves enabling the backbone LLM to perform reasoning over sophisticated prompts [wei2025univideo], and providing dense contextual embeddings to JAV-DiT that also requires fine-tuning to adapt to the shifted input distribution.
Discussion 2: the ultimate framework for joint audio-video comprehension and generation. The encoder–LLM–decoder architecture adopted in this work is a well-established and widely validated paradigm [zhan2024anygpt, wu2024next, lu2024unifiedio2], characterized by continuous token representations on the input side and a diffusion-based decoder/generator on the output side. In parallel, a growing number of unified models have emerged that perform end-to-end auto-regression (AR) over multimodal tokens [wang2024emu3, wu2024vila, team2024chameleon, tong2024metamorph]. Both discrete [wu2024vila] and continuous [tong2024metamorph] AR-based approaches have demonstrated significant mutual enhancement between understanding and generation tasks. Building upon these advances, extending from text–image to joint text–audio–video understanding and generation holds great potential for achieving even more remarkable capabilities.

### v2

@@ Abstract
This paper presents JavisGPT, the first unified multimodal large language model (MLLM) for joint audio-video (JAV) comprehension and generation. JavisGPT has a concise encoder-LLM-decoder architecture, which has a SyncFusion module for spatio-temporal audio-video fusion and synchrony-aware learnable queries to bridge a pretrained JAV-DiT generator. This design enables temporally coherent video-audio understanding and generation from multimodal instructions. We design an effective three-stage training pipeline consisting of multimodal pretraining, audio-video fine-tuning, and large-scale instruction-tuning, to progressively build multimodal comprehension and generation from existing vision-language models. For instruction tuning, we construct JavisInst-Omni, a high-quality instruction dataset with over 200K GPT-4o-curated audio-video-text dialogues that cover diverse and multi-level comprehension and generation scenarios. On JAV comprehension and generation benchmarks, our experiments show that JavisGPT outperforms existing MLLMs, particularly in complex and temporally synchronized settings.
@@ 1 Introduction
Joint understanding and generation of multiple modalities has recently sparked increasing interest [tang2023codi, wu2024towards, fei2024vitron, tang2024codi2, xu2025qwen25omni, microsoft2025phi4minitechnicalreportcompact]. Existing efforts generally focus on image-text frameworks [zhou2024transfusion, xie2024show, team2024chameleon, pan2025metaquery, chen2025januspro], or treat video and audio understanding and generation as separate modalities [zhan2024anygpt, wu2024next, lu2024unifiedio2].
@@
We introduce JavisGPT, a bespoke MLLM for high-quality, unified comprehension and generation in joint audio-video (JAV) tasks, which has clear application potential in avatar chatbots, sounding movie curation, and music-driven video analysis. JavisGPT is a holistic system that accepts as input separate audio and videos, synchronized sounding videos, and user text prompts. After comprehending and reasoning over these input signals, it produces either textual responses and/or synchronized sounding videos. Our architecture supports a wide range of JAV scenarios, including question-answering in comprehending sounding audios, generation of sounding audios following user instructions, upon in-context reasoning, and interleaved text-audio-video generation in multi-turn user dialogues (ref. Fig.˜1). As a comprehensive and state-of-the-art system, we highlight the design of JavisGPT below (with detailed discussions with related works provided in Sec.˜2).
Model architecture. JavisGPT uses an encoder-LLM-decoder architecture (ref. Fig.˜2), with Qwen2.5 [yang2024qwen25] as the LLM backbone. The visual encoder is inherited from Qwen2.5-VL [bai2025qwen25vl], and the audio encoder is based on BEATs [chen2022beats]. Audio and video features, along with user prompts and learnable JavisQuery tokens, are passed to the LLM. To enable fine-grained spatiotemporal alignment, we propose a dedicated SyncFusion module that fuses audio and video representations into synchronized SyncAV tokens for unified comprehension. At the output stage, the LLM generates textual responses along with JavisCond tokens, which encode contextual semantics and serve as conditioning inputs for a pretrained JAV-DiT generator [liu2025javisdit]. We choose JavisDiT for its generation quality and flexibility, and incorporate hierarchical JavisQueries to provide spatiotemporal priors, further enhancing synchronization in audio-video generation.
@@
We compare JavisGPT with existing JAV-like LLMs on various JAV benchmarks, where JavisGPT yields impressive performance across all scenarios. Moreover, for audio-video joint generation, JavisGPT also significantly outperforms existing methods, producing high-quality and diverse sounding videos. We reiterate our main points below:
@@ 3.1 Synchrony-aware Audio-Video Comprehension
Synchronized audio-video fusion. Unlike previous MLLMs that straightforwardly concatenate video and audio features together [cheng2024videollama2, sun2024videosalmonn] or simply interleave audio-video frames [tang2024videosalmonn2, xu2025qwen25omni], JavisGPT introduces a SyncFusion module $\psi^{av}$ to explicitly model the spatio-temporal synchrony in sounding videos. As illustrated in Fig.˜3, we first temporally align the audio and video streams by evenly dividing the audio representations to match the number of video frames, transforming the audio features from $\mathbf{A}\in\mathbb{R}^{T_{a}\times M\times C}$ to $\mathbf{A}^{\prime}\in\mathbb{R}^{T_{v}\times(r\times M)\times C}$, where $r=\lceil T_{a}/T_{v}\rceil$, with zero-paddings applied if needed. Then, a shared cross-attention module is employed to inject audio cues into the frame-wise video representations, followed by a two-layer MLP to generate SyncAV representations. The length of SyncAV representations matches that of the vanilla extracted video representations, ensuring compatibility with the MRoPE positional encoding used in the Qwen2.5-VL [bai2025qwen25vl]. Each resulting token $e_{i,j}^{t_{v}}\in\mathbb{R}^{C}$ encodes a localized audio-visual event occurring at the $i$-th row, $j$-th column visual patch of frame $t_{v}$. By integrating audio semantics into visual features, the SyncFusion module enhances representational expressiveness and achieves precise spatio-temporal synchronization. This is crucial for downstream tasks requiring fine-grained audio-video understanding.
@@ 3.2 Spatio-Temporally Synchronized Audio-Video Generation
Synchrony-enhanced audio-video generation. To further strengthen the spatio-temporal synchrony of generated sounding videos, we follow liu2025javisdit to additionally design $N^{\prime}$ learnable query tokens $Q^{s}$ to estimate spatio-temporal prior (ST-prior) embeddings $\hat{\mathbf{s}}$ for auxiliary conditions. As shown in Fig.˜4, the semantic condition $\hat{\mathbf{c}}$ a coarse description of the overall audio-visual event (e.g., ‘a car starts its engine and leaves the screen’), while the ST-prior condition $\hat{\mathbf{s}}$ captures fine-grained synchrony. Specifically, the spatial prior governs where events occur and disappear (e.g., ‘the car is in the top-left corner of screen’), whereas the temporal prior determines when they begin and end (e.g., ‘sound starts at 2s, exits at 7s, fades at 9s’). To compute $\hat{\mathbf{s}}$, we employ a separate MLP projector $\phi_{\theta}^{s}$ to map the final-layer representations of the prior query tokens into the ST-prior condition space, aligning them with the ST-prior encoder used in JavisDiT [liu2025javisdit]. Finally, we combine the alignment loss between the predicted ST-prior embeddings $\hat{\mathbf{s}}$ and the ground-truth ST-priors $\mathbf{s}$ with the denoising objective, jointly optimizing both to strengthen the audio-video synchrony:
@@ 4 Training and Datasets
Table 1: Summary of the training pipeline. ‘$M$-PT’ denotes multimodal pre-training, ‘AV-FT’ refers to audio-video fine-tuning, and ‘$M$-InstTune’ means multimodal instruction-tuning.
@@ 4.3 Instruction-tuning for Interleaved Comprehension and Generation
5 Contribution Statement
This work is the result of a large and comprehensive project that aims to provide the community with an open-source tool and platform for joint understanding and generation of sounding video, which is the first of its kind. While most of the individual network components are known in the literature, it takes significant trials and errors to find a feasible plan, and connect and align them to form a state-of-the-art and yet concise system. In particular, the proposed three-stage training pipeline has proven useful for gradually improving model performance. In addition, the proposed JavisInst-Omni dataset represents a very rich mixture of single- and multi-modality data, with the majority being single-turn and the rest multi-turn. We believe the architecture, training methods, and datasets offer a very competitive baseline and have significant value for future data-centric and model-centric studies.
@@ 6 Experiments
This section presents the experimental results to evaluate the effectiveness of our proposed JavisGPT. For a fair comparison, we train JavisGPT following the three-stage pipeline described in Sec.˜4, with each stage running for one epoch. We use zero-shot evaluation for downstream tasks including audio-video comprehension and generation using their official evaluation protocols. Additional implementation and training details can be found in Appendix˜B.
@@ 6.1 Comparison with the State of the Art
Multimodal comprehension. As shown in Tab.˜2, our method achieves performance on par with advanced models on video/audio understanding, demonstrating strong unimodal capabilities. It maintains visual comprehension strength of Qwen2.5-VL [bai2025qwen25vl] backbone (70.2 vs. 70.5 on Perception [patraucean2023perception]) and even slightly outperforms Qwen2.5-Omni [xu2025qwen25omni] on MVBench [li2024mvbench] (video) and TUT2017 [mesaros2017tut] (audio). For audio-video synchronized understanding, JavisGPT surpasses both Qwen2.5-Omni and VideoLLaMA2.1 [cheng2024videollama2] on AVQA [yang2022avqa] and MU-AVQA [liu2024musicavqa], achieving top results on both benchmarks. We attribute this to the spatio-temporal alignment capability of the proposed SyncFusion and our synchrony-aware training strategy. Notably, those results are obtained with fewer training samples (1.5M in total), highlighting its strong data efficiency.
@@
Interleaved comprehension and generation. As no existing benchmark directly measures joint understanding and generation of sounding videos, we fill this gap by carefully curating a new evaluation set comprising 100 multi-turn QA dialogues covering four types of interleaved understanding-generation tasks. We employ five unbiased volunteers to conduct human evaluation over several dimensions, covering both textual and audio-video outputs. Full setup and detailed analysis are provided in Sec.˜E.1 The average results are reported in Fig.˜6, where our method consistently outperforms the two competing methods with joint comprehension-generation capabilities by a large margin. Notably, we observe (1) NExT-GPT [wu2024next] frequently refuses to respond or generates noisy audio-video outputs, severely impacting instruction-following and generation quality; (2) UnifiedIO-2 [lu2024unifiedio2] is inferior to context reasoning and proactive thinking, failing to connect user history (e.g., preferences) with current instructions.
@@ 6.2 Further Analysis
Joint training improves both understanding and generation. Sec.˜6.2 compares the performance of comprehension and generation tasks under separate vs. joint training, with each score averaged over the corresponding test sets (e.g., ‘V-Und’ is the average accuracy over ActivityNet [caba2015activitynet], Perception [patraucean2023perception], and MVBench [li2024mvbench]). Results show that joint training consistently outperforms separate training in both understanding and generation tasks, indicating a mutual enhancement effect. The improvement is more substantial on the generation side.
To better understand this effect, Sec.˜6.2 presents the evolution of generation metrics across three dimensions. As the number of generation training samples increases, the inclusion of comprehension data leads to steady gains in both AV-Quality and Text-Consistency. It suggests that semantic understanding of AV content enhances the quality of condition embeddings used for generation. This trend also aligns with findings in MetaQueries [pan2025metaquery] for image tasks. However, AV synchrony shows only marginal improvement (+0.04), possibly indicating the need for a more unified mechanism bridging comprehension and generation. We further discuss this in Appendix˜A.
Case study: instruction-followed and synchrony-aware comprehension and generation. Fig.˜9 presents a qualitative comparison with UnifiedIO-2 [lu2024unifiedio2] and NExT-GPT [wu2024next] in a two-round dialogue setting. In the first round, the models are asked to identify the source of a gunfire sound. UnifiedIO-2 gives a general response (‘a semi-automatic weapon’), while NExT-GPT hallucinates an explosion of a flower. In contrast, our model accurately grounds the sound in a video game scene, showing better spatio-temporal grounding and scene understanding. In the second round, the models are asked to generate a gunfire-related video. UnifiedIO-2 produces irrelevant key frames, NExT-GPT fails to generate, while our JavisGPT successfully synthesizes coherent and synchronized audio-visual content. More examples are presented in Appendix˜F.
@@ 7 Conclusion and Discussion
We propose JavisGPT, a unified MLLM for synchronized audio–video comprehension and generation. JavisGPT adopts an encoder–LLM–decoder architecture with a SyncFusion module for spatiotemporal alignment and integrates a pretrained JavisDiT generator via hierarchical condition embedding. A three-stage training pipeline and the large-scale JavisInst-Omni dataset enable complex audio–video–text interactions. Extensive experiments demonstrate state-of-the-art performance on both unimodal and synchronized AV tasks, with interesting observations on mutual enhancement on both understanding and generation capabilities. We believe JavisGPT offers a strong foundation for future research in synchronized media generation.
Discussion 1: the extended capabilities for sounding video comprehension and generation. In the future, JavisGPT can be further extended to achieve higher fidelity in several directions: (1) Speech input and output by replacing the BEATs tokenizer with more advanced speech models (e.g., Whisper or WavTokenizer) [chen2022beats, radford2023whisper, ji2024wavtokenizer, xu2025qwen25omni]; (2) Fine-grained controllable audio–video generation and editing through additional conditioning signals in JAV-DiT [jiang2025vace]; (3) Complex instruction-based generation, requiring enhanced LLM reasoning and dense contextual conditioning for JAV-DiT [wei2025univideo].
Discussion 2: the ultimate framework for joint audio-video comprehension and generation. The encoder–LLM–decoder architecture is a well-established paradigm with continuous input tokens and diffusion-based decoders [zhan2024anygpt, wu2024next, lu2024unifiedio2]. In parallel, recent unified models adopt end-to-end autoregressive multimodal token modeling [wang2024emu3, wu2024vila, team2024chameleon, tong2024metamorph], where both discrete and continuous representations show strong synergy between understanding and generation. Extending these approaches from text–image to joint text–audio–video modeling offers substantial potential for further capability gains.